How many different ways can a committee of 3 be selected from among the 55 staff members at a bank?
Answer by jim_thompson5910(35256)

C(n,r) = (n!)/(r!(n-r)!)

C(55,3) = (55!)/(3!*(55-3)!)

C(55,3) = (55!)/(3!*52!)

C(55,3) = (55*54*53*52!)/(3!*52!)

C(55,3) = (55*54*53)/(3!)

C(55,3) = (55*54*53)/(3*2*1)

C(55,3) = (157410)/(6)

C(55,3) = 26235

So there are 26235 different ways to form the committee (assuming order doesn't matter; ie, each member of the committee is of the same rank)